OpenStudy (anonymous):

cos (theta-60)=3 cos theta find the general solution

OpenStudy (anonymous):

cos (theta-60)=3 cos theta cos theta cos 60 + sin theta son 60 = 3 cos theta as cos 60 = 1/2 and sin 60 = sqrt 3/2 (1/2) cos theta + sqrt 3 / 2 sin theta = 3 cos theta cos theta +sqrt 3 sin theta = 6 cos theta

OpenStudy (anonymous):

5 cos theta - sqrt 3 sin theta = 0 5 cos theta = sqrt 3 sin theta squaring both sides we get 25 cos ^2 theta = 3 sin^2 theta divide both sides by cos ^2 theta we get 25 = 3 (sin ^2 theta )/(cos^ theta) 25/3 = tan ^2 theta now solve for theta